OpenStudy (mven):

Find the surface area of a sphere that has a volume of 288 cu. in.

OpenStudy (mven):

144pi ?

OpenStudy (ilovecake):

A=4*pi*r^2 A=4*36*pi =144*pi

OpenStudy (ilovecake):

you are right @MVen

OpenStudy (mven):

awesome thank you

OpenStudy (ilovecake):

Your Welcome :)